# On solutions of singular Sylvester equations in quaternions

### 07 Mar 2026 v2

@@ 1. Introduction
Johnson [2] distinguished between the regular and singular cases and provided topological proofs for the existence and uniqueness of solutions over division rings, which directly apply to quaternions. A regular Sylvester equation has a unique solution for any quaternion c{c}. Hence, the homogeneous regular Sylvester equation has only the trivial solution x=0{x}=0. Closed-form expressions for the inhomogeneous regular Sylvester equation are given in [1, 3, 4].
The singular Sylvester equation over quaternions poses greater challenges. Tian provided solutions for both the homogeneous [5] and inhomogeneous [6] cases, but without explicit proofs, presenting results in forms that obscure the connection to the original problem. Shpakivsky [7] reduced the problem to a system of linear equations over the real number field, while Turner [8] separated real and imaginary components, also transforming the problem into the real domain. Both approaches are algebraically complex, obscuring the link between their structure and the original quaternionic formulation and hindering extension to related problems. Subsequent work, such as that by Janovská and Opfer [1], managed only to characterize the solution space and propose numerical algorithms rather than analytic, closed-form results.
We present both the general and nonzero solutions of the homogeneous singular Sylvester equation in quaternions, establishing the connection to the quaternionic roots. We also restate the solvability conditions with proofs and provide the general solution for the inhomogeneous singular Sylvester equation.
@@ 2. Quaternion Fundamentals
and satisfies the following properties
@@
In Definition 2.5, similarity was introducted only for nonreal quaternions. In general, however, this term may also be defined for real quaternions. For real similar quaternions a,b{a},{b} it holds ${a}=\mathrm{Re}\,\mathopen{(}{a}{p}{p}^{-1}\mathclose{)}=\mathrm{Re}\,\mathopen{(}{p}^{-1}{a}{p}\mathclose{)}={b}$ and further
@@
such that every nonzero quaternion p{p} possesses this property. This implies, that any quaternion is a solution to the homogeneous singular Sylvester equation, and that the inhomogeneous singular Sylvester equation has no solutions, as will be shown later.
@@ 3. Quaternion square roots
For any nonzero quaternion a{a} (3.1) has the solutions
@@
Now assume that a{a} is positive. If $\mathrm{Re}\,{x}{}=0$, then $a=\mathrm{Re}\,({x}{x})=-\mathopen{\lvert}{\mathrm{Im}\,{x}}\mathclose{\rvert}^{2}$, which contradicts that a{a} is a positive real number. Hence, we must have ${\mathrm{Im}\,{x}}=0$ and further ${x}=\mathrm{Re}\,{x}{}=\pm\textstyle\sqrt{a}$. Since ${a}=\mathopen{\lvert}{a}\mathclose{\rvert}$ it holds ${a}+\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\big\lvert}{a}+\mathopen{\lvert}{a}\mathclose{\rvert}\mathclose{\big\rvert}$, which yields (3.2). Conversely, assume that a{a} is negative. If ${\mathrm{Im}\,{x}}=0$, then ${a}=\mathrm{Re}\,({x}{x})=\mathopen{\lvert}\mathrm{Re}\,{x}{}\mathclose{\rvert}^{2}$, which contradicts that a{a} is a negative real number. Hence, we must have $\mathrm{Re}\,{x}{}=0$. Further, for any pure quaternion p{p} it holds ${p}{p}=-\mathopen{\lvert}{p}\mathclose{\rvert}^{2}$ such that
@@
The square roots of any nonzero quaternion a{a} are unit quaternions if and only if quaternion a{a} is unit.
@@
The square roots of any nonreal quaternion a{a} have nonzero real parts, i.e. the square roots of a nonreal quaternion a{a} are not pure.
@@
For any nonreal quaternion a{a} there exist nonzero real numbers $\lambda_{0},\lambda_{1}$ such that
@@
For any quaternions a,b{a},{b} with $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$, for which the quaternion product ab{a}{b} is nonreal, (3.5) has the solutions
@@ 4. Solutions to the homogeneous singular Sylvester equation
is referred to as a homogeneous singular Sylvester equation. In this section, we show the existence of solutions to (4.1) and derive their explicit representation. More precisely, we describe both the general and nontrivial solutions, where in the latter case, the quaternion square root plays a particularly significant role.
Theorem 4.1.
@@
where $\lambda=-4\mathopen{\lvert}{a}\mathclose{\rvert}^{2}$. Hence, any solution to (4.1) can be obtained from (4.3) with an appropriate q{q}. Now, let q{q} be an arbitrary quaternion. Then, by substituting (4.3) into (4.1) we obtain
@@
Hence, (4.3) is the general solution to (4.1) for any quaternion q{q}. ∎
@@
By Corollary 4.1.1, the quaternions a,b{a},{b} may, without loss of generality, be considered as pure. We examine the following cases to find the nonzero solutions to (4.1). Case 1: Let ${a}{b}\neq{b}{a}$. Then, by Lemma 2.1, it holds ${\mathrm{Im}\,{a}}\neq\nu({\mathrm{Im}\,{b}})$ for all real numbers $\nu$. Hence, any quaternion q{q} can be given as
@@
By applying Lemma 3.3 to $(\mathopen{\lvert}{a}\mathclose{\rvert}^{2}-{a}{b})$, and by Lemma 2.3, we further obtain (4.4). Case 2: Let ab=ba{a}{b}={b}{a}. Then, by Lemma 2.2 either a=b{a}={b} or a=b{a}=-{b} holds. Case 2a: Let a=b{a}={b}. Then, by Theorem 4.2, for any quaternion q{q}, the general solution to (4.1) is given as
@@
Let $\lambda,\mu$ be arbitrary real numbers. Then, with ${a}(\mathrm{Re}\,{q}{})=\mu\,({a}+{b})$ and ${a}\cdot{\mathrm{Im}\,{q}}=\lambda\sqrt{{a}{b}^{*}}$, according to Lemma 3.2, we obtain (4.4). Case 2b: Let a=b{a}=-{b}. Then, by Theorem 4.2, and for any quaternion q{q}, the general solution to (4.1) is given as
@@ 5. Solutions to the inhomogeneous singular Sylvester equation
Theorem 5.2.
Let a,b,c{a},{b},{c} be nonzero quaternions with ${a}\sim{b}$ and ${a}{c}={c}{b}^{*}$. Then, for any quaternion q{q}, the general solution to (5.1) is given as
@@
Applying Theorem 4.2 to the homogeneous singular Sylvester equation (5.3), we have
@@
for any quaternion q{q}. Substituting ${a}{c}=\frac{1}{2}({a}{c}-{c}{b})$ into (5.4) yields (5.2). ∎
Corollary 5.2.1.
@@
${x}=\lambda\sqrt{({\mathrm{Im}\,{a}})({\mathrm{Im}\,{b}}^{*})}+\mu\mathrm{Im}\,({a}+{b})+\frac{{c}}{4\mathopen{\lvert}{\mathrm{Im}\,{a}}\mathclose{\rvert}^{2}}\mathrm{Im}\,({b}-{a}).$

### 24 Mar 2026 v3

@@ 1. Introduction
Johnson [2] distinguished between the regular and singular cases and provided topological proofs for the existence and uniqueness of solutions over division rings, which directly apply to quaternions. A regular Sylvester equation has a unique solution for every quaternion c{c}. Hence, the homogeneous regular Sylvester equation has only the trivial solution x=0{x}=0. Closed-form expressions for the inhomogeneous regular Sylvester equation are given in [1, 3, 4].
The singular Sylvester equation over quaternions poses greater challenges. Rodman [5] provided a comprehensive matrix-theoretic foundation for quaternionic linear algebra, including detailed analyses of eigenvalues and similarity transformations, yet stopped short of providing explicit closed-form solutions for the singular Sylvester equation. Tian provided solutions for both the homogeneous [6] and inhomogeneous [7] cases, but without explicit proofs, presenting results in forms that obscure the connection to the original problem. Shpakivsky [8] reduced the problem to a system of linear equations over the real number field, while Turner [9] separated real and imaginary components, also transforming the problem into the real domain. Both approaches are algebraically complex, obscuring the link between their structure and the original quaternionic formulation and hindering extension to related problems. Subsequent work, such as that by Janovská and Opfer [1], managed only to characterize the solution space and propose numerical algorithms rather than analytic, closed-form results.
The main contribution of this work is the derivation of the analytic, closed-form nonzero solutions for the homogeneous and the general solution of the inhomogeneous singular Sylvester equations, as presented in Theorems 4.3 and 5.3, respectively. By establishing a structural connection to quaternionic square roots, we uncover a direct geometric decomposition of the solution space. This formulation facilitates a deeper understanding of the solution’s orientation relative to the axes of a{a} and b{b}, which is essential for engineering applications involving spatial rotations.
@@ 2. Quaternion Fundamentals
where the symbols (,×)($\cdot\,$$\times$) represent the vector scalar and cross product, respectively. The quaternion product satisfies the following properties
@@
In Definition 2.5, similarity was restricted to nonreal quaternions. However, the concept extends naturally to the real case. For real similar quaternions a,b{a},{b} it holds ${a}=\mathrm{Re}\,\mathopen{(}{a}{p}{p}^{-1}\mathclose{)}=\mathrm{Re}\,\mathopen{(}{p}^{-1}{a}{p}\mathclose{)}={b}$ and further
@@
such that every nonzero quaternion p{p} possesses this property. This implies that in the real case, every quaternion is a solution to the homogeneous singular Sylvester equation. Conversely, the inhomogeneous singular Sylvester equation has no solutions, as will be shown later.
@@
Remark 2.8.
The result concerning the similarity of the quaternions a{a} and b{b}, as stated in Lemma 2.7, is already present in the equivalence between (1) and (5) of Theorem 2.2.6 in [5]. However, the proof of this theorem does not provide a formula for explicitly determining a quaternion p{p} that establishes the similarity between a{a} and b{b}. In contrast, the proof of Lemma 2.7 explicitly constructs such a p{p} by considering cases 1, 2a, and 2b. This distinction is crucial as it enables the explicit determination of the nonzero solutions to the homogeneous singular Sylvester equation in Theorem 4.3.
@@ 3. Quaternion square roots
For an arbitrary nonzero quaternion a{a} (3.1) has the solutions
@@
Now assume that a{a} is positive. If $\mathrm{Re}\,{x}{}=0$, then $a=\mathrm{Re}\,({x}{x})=-\mathopen{\lvert}{\mathrm{Im}\,{x}}\mathclose{\rvert}^{2}$, which contradicts that a{a} is a positive real number. Hence, we must have ${\mathrm{Im}\,{x}}=0$ and further ${x}=\mathrm{Re}\,{x}{}=\pm\textstyle\sqrt{a}$. Since ${a}=\mathopen{\lvert}{a}\mathclose{\rvert}$ it holds ${a}+\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\big\lvert}{a}+\mathopen{\lvert}{a}\mathclose{\rvert}\mathclose{\big\rvert}$, which yields (3.2). Conversely, assume that a{a} is negative. If ${\mathrm{Im}\,{x}}=0$, then ${a}=\mathrm{Re}\,({x}{x})=\mathopen{\lvert}\mathrm{Re}\,{x}{}\mathclose{\rvert}^{2}$, which contradicts that a{a} is a negative real number. Hence, we must have $\mathrm{Re}\,{x}{}=0$. Further, for every pure quaternion p{p} it holds ${p}{p}=-\mathopen{\lvert}{p}\mathclose{\rvert}^{2}$ such that
@@
The square roots of an arbitrary nonzero quaternion a{a} are unit quaternions if and only if quaternion a{a} is unit.
@@
The square roots of an arbitrary nonreal quaternion a{a} have nonzero real parts, i.e. the square roots of a nonreal quaternion a{a} are not pure.
@@
For an arbitrary nonreal quaternion a{a} there exist nonzero real numbers $\lambda_{0},\lambda_{1}$ such that
@@
Remark 3.3.
A formula for calculating the quaternion square root is already provided in a similar form in Theorem 2.5.2 of [5]. However, the representation used in Lemma 3.2 is more suitable for practical applications in engineering systems, as the general form according to (3.2) applies to arbitrary quaternions a{a}. The only difference lies in the specific representation of quaternion q{q}, which depends on a{a}. In [5] however, completely different representations depending on a{a} were provided.
@@
For arbitrary quaternions a,b{a},{b} with $\mathopen{\lvert}{a}\mathclose{\rvert}=\mathopen{\lvert}{b}\mathclose{\rvert}$, for which the quaternion product ab{a}{b} is nonreal, (3.5) has the solutions
@@ 4. Solutions to the homogeneous singular Sylvester equation
is referred to as a homogeneous singular Sylvester equation. In this section, we restate the existance and the general solution to (4.1) previously presented in [6]. Moving beyond existing theory, we provide a novel characterization of the nontrivial solutions to (4.1) by applying the quaternion square root, presented in Theorem 4.3.
Lemma 4.1.
@@
where $\lambda=-4\mathopen{\lvert}{a}\mathclose{\rvert}^{2}$. Hence, every solution to (4.1) can be obtained from (4.3) with an appropriate q{q}. Now, let q{q} be an arbitrary quaternion. Then, by substituting (4.3) into (4.1), we obtain
@@
Hence, (4.3) is the general solution to (4.1) for an arbitrary quaternion q{q}. ∎
@@
By Corollary 4.1.1, the quaternions a,b{a},{b} may, without loss of generality, be considered as pure. We examine the following cases to find the nonzero solutions to (4.1). Case 1: Let ${a}{b}\neq{b}{a}$. Then, by Lemma 2.1, it holds ${\mathrm{Im}\,{a}}\neq\nu({\mathrm{Im}\,{b}})$ for all real numbers $\nu$. Hence, every quaternion q{q} can be given as
@@
By applying Lemma 3.4 to $(\mathopen{\lvert}{a}\mathclose{\rvert}^{2}-{a}{b})$, and by Lemma 2.3, we further obtain (4.4). Case 2: Let ab=ba{a}{b}={b}{a}. Then, by Lemma 2.2 either a=b{a}={b} or a=b{a}=-{b} holds. Case 2a: Let a=b{a}={b}. Then, by Lemma 4.2, for an arbitrary quaternion q{q}, the general solution to (4.1) is given as
@@
Let $\lambda,\mu$ be arbitrary real numbers. Then, with ${a}(\mathrm{Re}\,{q}{})=\mu\,({a}+{b})$ and ${a}\cdot{\mathrm{Im}\,{q}}=\lambda\sqrt{{a}{b}^{*}}$, according to Lemma 3.2, we obtain (4.4). Case 2b: Let a=b{a}=-{b}. Then, by Lemma 4.2, for an arbitrary quaternion q{q}, the general solution to (4.1) is given as
@@
Remark 4.4.
The representation in (4.4) identifies the quaternionic directions that span the solution space of (4.1). Specifically, the solution is decomposed into two orthogonal components aligned with the sum $\mathrm{Im}\,({a}+{b})$ and the quaternion square root $\sqrt{({\mathrm{Im}\,{a}})({\mathrm{Im}\,{b}}^{*})}$. This explicit spanning set provides a direct geometric link between the equation parameters and the solution’s orientation. Such structural transparency is vital in engineering systems, where the physical interpretation of quaternionic variables relative to rotation axes is critical.
@@ 5. Solutions to the inhomogeneous singular Sylvester equation
Lemma 5.2.
Let a,b,c{a},{b},{c} be nonzero quaternions with ${a}\sim{b}$ and ${a}{c}={c}{b}^{*}$. Then, for an arbitrary quaternion q{q}, the general solution to (5.1) is given as
@@
Applying Lemma 4.2 to the homogeneous singular Sylvester equation (5.3), we have
@@
for an arbitrary quaternion q{q}. Substituting ${a}{c}=\frac{1}{2}({a}{c}-{c}{b})$ into (5.4) yields (5.2). ∎
Theorem 5.3.
@@
(5.5) ${x}=\lambda\sqrt{({\mathrm{Im}\,{a}})({\mathrm{Im}\,{b}}^{*})}+\mu\mathrm{Im}\,({a}+{b})+\frac{{c}}{4\mathopen{\lvert}{\mathrm{Im}\,{a}}\mathclose{\rvert}^{2}}\mathrm{Im}\,({b}-{a}).$
Proof.
From Lemma 4.2 and Theorem 4.3 for an arbitrary quaternion q{q} and arbitrary real numbers $\lambda,\mu$ it holds
$({\mathrm{Im}\,{a}}){q}+{q}({\mathrm{Im}\,{b}})=\lambda\sqrt{({\mathrm{Im}\,{a}})({\mathrm{Im}\,{b}}^{*})}+\mu\mathrm{Im}\,({a}+{b}),$
which applied to (5.2) yields (5.5). ∎
@@ 6. Conclusion
Data Availability Statement
Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.